OpenStudy (anonymous):

solve & check. medal and fan x^2-3x+5=x(x-6)+20

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{x^2-3x+5=x(x-6)+20}\) \(\normalsize\color{blue}{ x^2-3x+5=x^2-6x+20}\) \(\normalsize\color{blue}{ 0=-3x+15}\) \(\normalsize\color{blue}{ x=5}\)

OpenStudy (anonymous):

how do i check?